OpenStudy (anonymous):

Assuming x ≠ 0 and y ≠ 0, what is the quotient of ...

OpenStudy (anonymous):

\[24x^6y^2 + 16x^4y^3 + 8x^2y^4 \over 4x^2y\]

OpenStudy (anonymous):

@misty1212

OpenStudy (freckles):

separate the three fractions

OpenStudy (anonymous):

How? Can you walk me through it please? :)

OpenStudy (freckles):

\[\frac{24 x^6y^2}{4x^2y}+\frac{16x^4y^3}{4x^2y}+\frac{8x^2y^4}{4x^2y}\]

OpenStudy (freckles):

now simplify wach fraction

OpenStudy (freckles):

24/4=? x^6/x^2=? use quotient rule y^2/y=?

OpenStudy (anonymous):

Okay, gimme a sec to do this on paper. Please don't leave yet.

OpenStudy (anonymous):

Would we have 6, x^4 and y?

OpenStudy (freckles):

yep

OpenStudy (anonymous):

Okay. Do you know what to do from here?

OpenStudy (freckles):

\[\frac{24 x^6y^2}{4x^2y}+\frac{16x^4y^3}{4x^2y}+\frac{8x^2y^4}{4x^2y} \\ 6x^4y+\frac{16x^4y^3}{4x^2y}+\frac{8x^2y^4}{4x^2y} \]

OpenStudy (freckles):

you need to do the other two fractions

OpenStudy (anonymous):

Also, I apologize for any delayed replies. My computer is having issues.

OpenStudy (anonymous):

Okay. One minute.

OpenStudy (anonymous):

Sorry I had something burning in the kitchen.

OpenStudy (anonymous):

Would we have 4, x^2 and y^2?

OpenStudy (anonymous):

For the second fraction

OpenStudy (anonymous):

And for the third fraction would we have 2, x and y^3?

OpenStudy (freckles):

the third fraction is just a lil bit off

OpenStudy (freckles):

anything not equal to 0 where you have that anything divide by itself is 1

OpenStudy (freckles):

that is like example 5/5=1 6/6=1 -5/-5=1 x not 0 x/x=1 x^2/x^2=1

OpenStudy (anonymous):

I thought that a variable always had ^1 but it was not seen. For example, instead of x^1 it would be x.

OpenStudy (freckles):

x^1 is x

OpenStudy (freckles):

x^0 is 1

OpenStudy (anonymous):

Oh I understand now! Thank you! :)

OpenStudy (freckles):

np